# Lensing and wave optics in the strong field of a black hole

### v2

@@ IV.1 Lensed Binary Black Hole System and Wave Optics Effects
To illustrate the effects of lensing on the GW signal, we simulate a GW waveform from the merger of two non-spinning BHs with equal masses (in the detector frame) $30\leavevmode\nobreak\ M_{\odot}$, a representative case for current ground-based detectors. We adopt a typical redshift $z=0.1$, which lies within the redshift horizon of GW detectors in the third LIGO-Virgo-KAGRA (LVK) observing run [78], corresponding to a luminosity distance of $475.8\leavevmode\nobreak\ \rm{Mpc}$. Inclination $\iota$ is set to zero, resulting in equal magnitude of the $+$ and $\times$ polarizations. We use the IMRPhenomXPHM [79] waveform approximant in LALSimulation [80]. A detailed summary of the waveform parameters is listed in Table 1, with definitions following [80].
@@
The lens parameters are chosen to ensure the gravitational wave (GW) signal remains in the wave optics regime for the binary black hole (BH) system under study. With a peak frequency of $f_{\rm peak}\sim 200\mathrm{Hz}$, a lens mass in the range $M\in[100,1000]M_{\odot}$ satisfies $62\pi f_{\rm peak}(1+z)M\sim 0.6-$ 6, placing us in the intermediate regime between deep wave optics and geometric optics—our primary region of interest. This motivates our choice of lens mass within $M\in[100,1000]M_{\odot}$. Given the comparable masses of the lens and merging BHs, the system likely forms a triple. If $r_{\rm LS}\omega\gg 1$, the amplification factor approaches $\mathcal{F}\to 1$ and $\mathcal{G}\to 0$, making lensing effects negligible. To avoid this regime, we set $r_{\rm LS}=100M$, which, while close to the merging BHs, remains sufficient to justify the plane wave approximation. The scales considered in this problem are comparable to those of [14].
@@
We emphasize that when relating distances, it is important to account for cosmic expansion by converting physical distances into angular diameter distances. Although this effect is not critical here due to the moderate redshift, $z=0.1$, properly incorporating cosmological distances ensures that the strong-field lensing formalism remains valid for arbitrarily distant lenses and sources. Details of this conversion are provided in App. B.
@@ IV.2 Frequency domain waveform
We begin by analyzing the lensed waveform in the frequency domain, shown in Fig. 5. As explained before, when accounting for the strong field effects, the two polarization modes, $+$ and $\times$ are lensed differently. This effect, which is a direct consequence of the tensorial nature of GWs, is unaccounted for in the PL approximation. Despite having a large impact parameter, we observe noticeable lensing-induced distortions to the GW. These are more evident for the strong field, BH lens , $M=100M_{\odot}$ (upper panel). Qualitatively, the PL approximation seems to capture the most prominent features of the lensed waveform.
In the high mass case, $M=1000M_{\odot}$ (lower panel), we observe that while the PL approximation predicts a highly oscillatory strain at all frequencies, the strong field calculation shows the amplitude of these oscillations decreasing with increasing frequency. This is the same feature shown in Fig. 4 at high frequencies, GWs of the lensed contribution are absorbed by the BH, and we recover only the unlensed component.
Figure 5: Frequency domain GW waveform for a lens mass $M=100M_{\odot}$ (top), and $M=1000M_{\odot}$ (bottom). The unlensed waveform is shown as a gray wide line, and it has equal magnitude for both $+$ and $\times$ polarizations. Under the PL approximation, the polarization content is not affected. The lensed waveform is shown as a red line. Finally, when accounting for the strong field effects, both polarizations are affected differently, as shown in blue ($+$), and green ($\times$).
@@
In the bottom panel, we consider the lens mass $M=1000M_{\odot}$, which begins to probe the regime where $\omega M\gtrsim 1$. Here, BH absorption becomes more significant, and higher frequencies are not lensed in our strong-field calculation. This is partly due to an assumption in our model that the transmitted component is unaffected by the gravitational field of the BH — a reasonable assumption in the wave optics regime, but less valid as the system approaches the geometric optics regime. In the latter, GWs propagate along null geodesics of the metric. Further work is needed to establish a robust connection between the wave optics and geometric optics regimes in the context of GW lensing by BHs.
@@
Figure 6: Absolute value of the difference between the lensed and unlensed waveforms when lensed by a BH taking into account the effects of the strong gravitational field, for the $+$ (blue) and $\times$ (green), and using the point mass lens approximation (red). Recall that the PL approximation lenses equally both polarizations. The top (bottom) panel shows the case where the lens mass is $M=100M_{\odot}$ (respectively, $M=1000M_{\odot}$).
Figure 7: Time domain GW waveform observed at the detector aligned at peak, unlensed (black), lensed by a BH, accounting for strong field effects (blue), and in the weak field, PL approximation (red). The left (respectively, right) panel corresponds to a lens mass $M=100M_{\odot}$ (resp. $M=1000M_{\odot}$). In order to obtain the detector strain, we assume a sky location given by the right ascension to be $\pi/8$, declination to be $\pi/4$ and polarization angle to be $\pi/4$.
@@ IV.3 Time domain waveform
Fig. 7 presents the time-domain waveforms for both lens masses previously considered. We display three cases—unlensed, lensed by a BH in the strong-field regime, and lensed using the weak-field PL approximation—aligning them by their peak amplitudes. Due to the large impact parameter, the lensed and unlensed signals exhibit a very small dephasing. However, magnification effects are apparent. For the low mass case, $M=100M_{\odot}$, the PL approximation predicts a slight demagnification, while the strong field calculation reveals a subtle magnification, which becomes most relevant close to the merger. The situation is different for the high mass scenario, $M=1000M_{\odot}$. Both methods show magnification, which is stronger in the BH scenario. Notably, at higher frequencies, black hole absorption reduces distortions near the merger in the strong-field calculation—an effect absent in the PL approximation. On the other hand, the distortions to the waveform near the merger are very relevant for the PL approximation.
We further analyze distortions during the inspiral phase for the more massive lens. Fig. 8 focuses on the early inspiral for the unlensed, BH-lensed, and PL-lensed waveforms. Despite the large impact parameter, the PL approximation exhibits a subtle beating pattern, arising from the convolution of the waveform with the high-frequency oscillations of the amplification factor (see Fig. 4). In contrast, the strong-field amplification factor rapidly approaches $|\mathcal{F}|\to 1$ as $\omega M\gg 1$, suppressing such interference effects. For smaller impact parameters, we expect more pronounced beating patterns in both regimes [19, 14]. These patterns encode information about the absorption of high-frequency radiation by the black hole, and resolving their structure will be the focus of future work.
@@ IV.4 Mismatch
We quantify the differences in the results between taking the full strong field nature of the BH into account as opposed to making the point mass lens approximation. By changing the mass of the lens $M\in[100,1000]M_{\odot}$ while keeping the GW being lensed fixed, we can scan different regimes -from the pure wave optics regime to the transition towards geometric optics. In order to quantify the distortions introduced to the GW by gravitational lensing, we define the mismatch as
@@
Figure 8: GW strain at the detector for the original waveform (black, left), the waveform lensed by a BH with mass $M=1000M_{\odot}$ (blue, middle), and lensed by a PL with the same mass, in the weak field regime (red, right). A beating pattern is evident in this last case.
Figure 9: Mismatch $\epsilon$ between lensed and unlensed waveforms defined in Eq. (37) as a function of the lens mass, for the PL approximation (red), and for the BH lens ($+$ and $\times$ polarizations in blue, and green, respectively).
Fig. 9 shows the mismatch between lensed and unlensed waveform both for the BH and the point mass lenses, as a function of the lens mass. The mismatch is significantly different between both methods considered here. Although part of this difference is due to the polarization mixing contributions, the dependence of the mismatch on the lens mass is also different. We find that $\epsilon$ becomes smaller for larger values of the lens mass. This is due to a larger fraction of the lensed component of the waveform being absorbed by the BH. This is consistent with the findings of [39], where they also report a better agreement between the lensed waveform and the reconstructed, best-fit waveform, for larger lens masses, compared to smaller lens masses. Due to the large impact parameter considered here, we only find a small mismatch, $\epsilon\lesssim 1\%$, whereas [39] considers the perfectly aligned case. However, the methods presented here can be improved upon in order to understand the dependence of the mismatch not only on the lens mass, but also on the impact parameter, in the strong field regime.
@@ V Discussion
In this work, we explored the gravitational lensing of GWs by the strong gravitational field of a Schwarzschild BH. While this problem has a long history, a direct comparison between wave scattering in BH perturbation theory and the weak-field approximation in GW lensing was lacking. We present, for the first time, a semi-analytical study of GW lensing in the strong-field, wave-optics regime. Previous works had either focused on the deep wave-optics limit ($\omega M\ll 1$[40], considered only wave scattering [41, 42], or relied on fully numerical time-domain analyses [39].
@@
We began by reviewing the point mass lens approximation in the weak field regime, and compared it with the equivalent calculation in the strong field. This involves decomposing the incoming plane wave in spherical harmonics, propagating those through the curved geometry, where they are modified through the BH reflection coefficients, and finally reconstructing the plane wave by resumming the multipolar expansion. To ensure convergence of the resummation, we introduced an averaging procedure dubbed Cesàro summation. Despite this, convergence remains slow at smaller inclination angles, limiting our ability to explore small impact parameters. This could be improved by incorporating, e.g., semi-analytical approximations for the reflectivity coefficients for large $\ell$ modes, though further exploration is needed. Using this resummation, we derived a direct expression for the GW amplification factor in the strong-field regime, parametrized by the BH’s reflection coefficients. This framework is flexible and can be extended to study lensing by other compact objects, including neutron stars, exotic objects, or BHs in modified gravity theories. A restricting assumption of our work is that we decompose the GWs in a directly transmitted component, which is unaffected by the BH potential, and a lensed component which is distorted following [40]. In the geometric optics limit, the lensed component is almost fully absorbed by the BH, and we only recovered the unlensed waveform. Relaxing this assumption will allow for a more clear connection between the wave and geometric optics limits, in the strong field regime.
@@
We applied our formalism to a simple GW signal (not unlike the first observed LVK event), considering an intermediate-mass BH lens ($M=10^{2}-10^{3}M_{\odot}$) in the $\omega M\sim 1$ regime. The lensed waveform is magnified, with a frequency dependence which shows interference effects. For $\omega M\lesssim 1$ the point mass lens approximation captures accurately the main features of gravitational lensing by a BH. We find disagreements mostly stemming from the polarization mixing contributions, which are neglected in the weak field regime, but important when the lens is a compact object such as a BH. For $\omega M\gtrsim 1$, both methods disagree, as e.g. GW absorption by the BH becomes significant. This qualitative agreement is remarkable and should not be overlooked. It suggests that the point mass lens approximation could be systematically improved to account for absorption (or generic finite size effects) and polarization-mixing, offering a practical alternative to full relativistic treatments.
@@
In the time domain, we identify significant magnification effects due to gravitational lensing, which are most important near merger. The fully relativistic lensed signal shows some differences with respect to the point lens approximation, which need to be scrutinized further. In particular, when the lens mass is relatively high, the point mass lens approximation predicts the appearance of beating patterns during the inspiral. These are absent for the strongly gravitating, BH lens -the lensed component is mostly absorbed by the BH at large values of $\omega M$. Due to the large impact parameter that we consider, we do not observe a significant phase shift. However, significant dephasing between lensed and unlensed waveforms is expected for smaller impact parameters (smaller values of the angle $\theta_{S}$), also in the strong field regime. Finally, we show the behavior of the mismatch between lensed and unlensed components as a function of the lens mass. While the magnitude of the mismatch does not vary significantly for the point mass lens case, oscillating around $\epsilon\sim 0.7\%$, the mismatch evolves significantly for the strongly gravitating lens. We find, consistently with the result of [39] that the mismatch decreases with lens mass.
@@
Our results bridge the gap between prior strong-field lensing studies [39, 42, 40], which either restricted the observer to be positioned in the vicinity of the lens or relied on low-frequency expansions. However, our framework cannot yet describe observers exactly aligned with the lens and source ($\theta_{\rm OL}=0$), where maximum magnification is expected to occur. We also find no arrival-time difference between $|h_{+}|$ and $|h_{\times}|$, consistent with previous findings [8] for non-rotating BHs. This is not the case for a rotating BH lens.
@@ Acknowledgements
We thank Jose Ezquiaga, Vitor Cardoso, and Miguel Zumalacárregui for valuable discussions and insightful comments. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184 This work makes use of the Black Hole Perturbation Toolkit [72]. We thank the The Erwin Schrödinger International Institute for Mathematics and Physics (ESI) and the organizers of the workshop “Lensing and Wave Optics in Strong Gravity”. We acknowledge support by VILLUM Foundation (grant no. VIL37766 and no. VIL53101) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for supporting this work.

### v3

@@ IV.1 Lensed binary black hole system and wave optics effects
To illustrate the effects of lensing on the GW signal, we simulate a GW waveform from the merger of two nonspinning BHs with equal masses (in the detector frame) $30\leavevmode\nobreak\ M_{\odot}$, a representative case for current ground-based detectors. We adopt a typical redshift $z=0.1$, which lies within the redshift horizon of GW detectors in the third LIGO-Virgo-KAGRA (LVK) observing run [78], corresponding to a luminosity distance of $475.8\leavevmode\nobreak\ \rm{Mpc}$. Inclination $\iota$ is set to zero, resulting in equal magnitude of the ++ and $\times$ polarizations. We use the IMRPhenomXPHM [79] waveform approximant in LALSimulation [80]. A detailed summary of the waveform parameters is listed in Table 1, with definitions following [80].
@@
The lens parameters are chosen to ensure the GW signal remains in the wave optics regime for the BH system under study. With a peak frequency of $f_{\rm peak}\sim 200\mathrm{Hz}$, a lens mass in the range $M\in[100,600]M_{\odot}$ satisfies $2\pi f_{\rm peak}(1+z)M\sim 0.6-3.6$, placing us in the intermediate regime between deep wave optics and geometric optics—our primary region of interest. This motivates our choice of lens mass within $M\in[100,600]M_{\odot}$. Given the comparable masses of the lens and merging BHs, the system likely forms a triple. If $r_{\rm LS}\omega\gg 1$, the amplification factor approaches $\mathcal{F}\to 1$ and $\mathcal{G}\to 0$, making lensing effects negligible. To avoid this regime, we set $r_{\rm LS}=100M$, which, while close to the merging BHs, remains sufficient to justify the plane wave approximation. The scales considered in this problem are comparable to those of [14].
@@
We emphasize that when relating distances, it is important to account for cosmic expansion by converting physical distances into angular diameter distances. Although this effect is not critical here due to the moderate redshift, $z=0.1$, properly incorporating cosmological distances ensures that the strong-field lensing formalism remains valid for arbitrarily distant lenses and sources. Details of this conversion are provided in Appendix B.
@@ IV.2 Frequency domain waveform
We begin by analyzing the lensed waveform in the frequency domain, shown in Fig. 5. As explained before, when accounting for the strong field effects, the two polarization modes, ++ and $\times$ are lensed differently. This effect, which is a direct consequence of the tensorial nature of GWs, is unaccounted for in the PL approximation. Despite having a large impact parameter, we observe noticeable lensing-induced distortions to the GW. These are more evident for the strong field, BH lens , $M=100M_{\odot}$ (upper panel). Qualitatively, the PL approximation seems to capture the most prominent features of the lensed waveform.
In the high mass case, $M=200M_{\odot}$ (lower panel), we observe that while the PL approximation predicts a highly oscillatory strain at all frequencies, the strong field calculation shows the amplitude of these oscillations decreasing with increasing frequency. This is the same feature shown in Fig. 4 at high frequencies, GWs of the lensed contribution are absorbed by the BH, and we recover only the unlensed component.
Figure 5: Frequency domain GW waveform for a lens mass $M=100M_{\odot}$ (top), and $M=200M_{\odot}$ (bottom). The unlensed waveform is shown as a gray wide line, and it has equal magnitude for both ++ and $\times$ polarizations. Under the PL approximation, the polarization content is not affected. The lensed waveform is shown as a red line. Finally, when accounting for the strong field effects, both polarizations are affected differently, as shown in blue (++), and green ($\times$).
@@
In the bottom panel, we consider the lens mass $M=200M_{\odot}$, which begins to probe the regime where $\omega M\gtrsim 1$. Here, BH absorption becomes more significant, and higher frequencies are not lensed in our strong-field calculation. This is partly due to an assumption in our model that the transmitted component is unaffected by the gravitational field of the BH—a reasonable assumption in the wave optics regime, but less valid as the system approaches the geometric optics regime. In the latter, GWs propagate along null geodesics of the metric. Further work is needed to establish a robust connection between the wave optics and geometric optics regimes in the context of GW lensing by BHs.
@@
Figure 6: Absolute value of the difference between the lensed and unlensed waveforms when lensed by a BH taking into account the effects of the strong gravitational field, for the ++ (blue) and $\times$ (green), and using the point mass lens approximation (red). Recall that the PL approximation lenses equally both polarizations. The top (bottom) panel shows the case where the lens mass is $M=100M_{\odot}$ (respectively, $M=200M_{\odot}$).
Figure 7: Time domain GW waveform observed at the detector aligned at peak, unlensed (black), lensed by a BH, accounting for strong field effects (blue), and in the weak field, PL approximation (red). The left (respectively, right) panel corresponds to a lens mass $M=100M_{\odot}$ (resp. $M=200M_{\odot}$). In order to obtain the detector strain, we assume a sky location given by the right ascension to be $\pi/8$, declination to be $\pi/4$ and polarization angle to be $\pi/4$.
@@ IV.3 Time domain waveform
Figure 7 presents the time-domain waveforms for both lens masses previously considered. We display three cases—unlensed, lensed by a BH in the strong-field regime, and lensed using the weak-field PL approximation—aligning them by their peak amplitudes. Due to the large impact parameter, the lensed and unlensed signals exhibit a very small dephasing. However, magnification effects are apparent. For the low mass case, $M=100M_{\odot}$, the PL approximation predicts a slight demagnification, while the strong field calculation reveals a subtle magnification, which becomes most relevant close to the merger. The situation is different for the high mass scenario, $M=200M_{\odot}$. Both methods show magnification, which is stronger in the BH scenario. Notably, at higher frequencies, black hole absorption reduces distortions near the merger in the strong-field calculation—an effect absent in the PL approximation. On the other hand, the distortions to the waveform near the merger are very relevant for the PL approximation. We also observe the presence of the second image after the merger in both the PL approximation, with an increasing relative time delay and a similar relative magnification ($\sim 0.1$), indicating the contribution of geometrical optics limit in this regime.
We further analyze distortions during the inspiral phase for the more massive lens. Figure 8 focuses on the early inspiral for the unlensed, BH-lensed, and PL-lensed waveforms. Despite the large impact parameter, the PL approximation exhibits a subtle beating pattern, arising from the convolution of the waveform with the high-frequency oscillations of the amplification factor (see Fig. 4). In contrast, the strong-field amplification factor rapidly approaches $|\mathcal{F}|\to 1$ as $\omega M\gg 1$, suppressing such interference effects. For smaller impact parameters, we expect more pronounced beating patterns in both regimes [19, 14]. These patterns encode information about the absorption of high-frequency radiation by the black hole, and resolving their structure will be the focus of future work.
@@ IV.4 Mismatch
We quantify the differences in the results between taking the full strong field nature of the BH into account as opposed to making the point mass lens approximation. By changing the mass of the lens $M\in[100,600]M_{\odot}$ while keeping the GW being lensed fixed, we can scan different regimes–from the pure wave optics regime to the transition toward geometric optics. Small modulations of the waveform appear in this last case. In order to quantify the distortions introduced to the GW by gravitational lensing, we define the mismatch as
@@
Figure 8: GW strain at the detector for the original waveform (black, left), the waveform lensed by a BH with mass $M=200M_{\odot}$ (blue, middle), and lensed by a PL with the same mass, in the weak field regime (red, right). Small modulations of the waveform appear in this last case.
Figure 9: Mismatch $\epsilon$ between lensed and unlensed waveforms defined in Eq. (37) as a function of the lens mass, for the PL approximation (red), and for the BH lens (++ and $\times$ polarizations in blue, and green, respectively).
Figure 9 shows the mismatch between lensed and unlensed waveforms both for the BH and the point mass lenses, as a function of the lens mass. The mismatch is significantly different between both methods considered here. Although part of this difference is due to the polarization mixing contributions, the dependence of the mismatch on the lens mass is also different. We find that $\epsilon$ becomes smaller for larger values of the lens mass. This is due to a larger fraction of the lensed component of the waveform being absorbed by the BH. This is consistent with the findings of [39], where they also report a better agreement between the lensed waveform and the reconstructed, best-fit waveform, for larger lens masses, compared to smaller lens masses. Due to the large impact parameter considered here, we only find a small mismatch, $\epsilon\lesssim 1\%$, whereas [39] considers the perfectly aligned case. However, the methods presented here can be improved upon in order to understand the dependence of the mismatch not only on the lens mass, but also on the impact parameter, in the strong field regime.
@@ V Discussion
In this work, we explored the gravitational lensing of GWs by the strong gravitational field of a Schwarzschild BH. While this problem has a long history, a direct comparison between wave scattering in BH perturbation theory and the weak-field approximation in GW lensing was lacking. We present, for the first time, a semianalytical study of GW lensing in the strong-field, wave-optics regime. Previous works had either focused on the deep wave-optics limit ($\omega M\ll 1$[40], considered only wave scattering [41, 42], or relied on fully numerical time-domain analyses [39].
@@
We began by reviewing the point mass lens approximation in the weak field regime, and compared it with the equivalent calculation in the strong field. This involves decomposing the incoming plane wave in spherical harmonics, propagating those through the curved geometry, where they are modified through the BH reflection coefficients, and finally reconstructing the plane wave by resumming the multipolar expansion. To ensure convergence of the resummation, we introduced an averaging procedure dubbed Cesàro summation. Despite this, convergence remains slow at smaller inclination angles, limiting our ability to explore small impact parameters. This could be improved by incorporating, e.g., semianalytical approximations for the reflectivity coefficients for large $\ell$ modes, though further exploration is needed. Using this resummation, we derived a direct expression for the GW amplification factor in the strong-field regime, parametrized by the BH’s reflection coefficients. This framework is flexible and can be extended to study lensing by other compact objects, including neutron stars, exotic objects, or BHs in modified gravity theories. A restricting assumption of our work is that we decompose the GWs in a directly transmitted component, which is unaffected by the BH potential, and a lensed component which is distorted following [40]. In the geometric optics limit, the lensed component is almost fully absorbed by the BH, and we only recovered the unlensed waveform. Relaxing this assumption will allow for a more clear connection between the wave and geometric optics limits, in the strong field regime.
@@
We applied our formalism to a simple GW signal (not unlike the first observed LVK event), considering an intermediate-mass BH lens ($M=100-600M_{\odot}$) in the $\omega M\sim 1$ regime. The lensed waveform is magnified, with a frequency dependence that shows interference effects. For $\omega M\lesssim 1$ the point mass lens approximation captures accurately the main features of gravitational lensing by a BH. We find disagreements mostly stemming from the polarization mixing contributions, which are neglected in the weak field regime, but important when the lens is a compact object such as a BH. For $\omega M\gtrsim 1$, both methods disagree, as e.g. GW absorption by the BH becomes significant. This qualitative agreement is remarkable and should not be overlooked. It suggests that the point mass lens approximation could be systematically improved to account for absorption (or generic finite size effects) and polarization mixing, offering a practical alternative to full relativistic treatments.
@@
In the time domain, we identify significant magnification effects due to gravitational lensing, which are most important near merger. The fully relativistic lensed signal shows some differences with respect to the point lens approximation, which need to be scrutinized further. In particular, when the lens mass is relatively high, the point mass lens approximation predicts the appearance of beating patterns during the inspiral. These are absent for the strongly gravitating, BH lens–the lensed component is mostly absorbed by the BH at large values of $\omega M$. Due to the large impact parameter that we consider, we do not observe a significant phase shift. However, significant dephasing between lensed and unlensed waveforms is expected for smaller impact parameters (smaller values of the angle $\theta_{S}$), also in the strong field regime. Finally, we show the behavior of the mismatch between lensed and unlensed components as a function of the lens mass. While the magnitude of the mismatch does not vary significantly for the point mass lens case, oscillating around $\epsilon\sim 0.7\%$, the mismatch evolves significantly for the strongly gravitating lens. We find, consistent with the result of [39] that the mismatch decreases with lens mass.
@@
Our results bridge the gap between prior strong-field lensing studies [39, 42, 40], which either restricted the observer to be positioned in the vicinity of the lens or relied on low-frequency expansions. However, our framework cannot yet describe observers exactly aligned with the lens and source ($\theta_{\rm OL}=0$), where maximum magnification is expected to occur. We also find no arrival-time difference between $|h_{+}|$ and $|h_{\times}|$, consistent with previous findings [8] for nonrotating BHs. This is not the case for a rotating BH lens.
@@ Acknowledgements
We thank Jose Ezquiaga, Vitor Cardoso, and Miguel Zumalacárregui for valuable discussions and insightful comments. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. This work makes use of the Black Hole Perturbation Toolkit [72]. We thank the The Erwin Schrödinger International Institute for Mathematics and Physics (ESI) and the organizers of the workshop “Lensing and Wave Optics in Strong Gravity”. We acknowledge support by VILLUM Foundation (grant no. VIL37766 and no. VIL53101) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for supporting this work.